Blue Water Systems is analyzing a project with the following cash flows. Should this project be accepted based on the modified internal rate of return if the discount rate is 12 percent? Why or why not?

Year   Cash flow

1    -260,000

2 85,000

3 128,600

4    96,780
Answer #1

Solution:-

To Calculate MIRR of the Project-

MIRR = \frac{Future Value of Cash Flows}{Present Value of Cash Outflow}^{\frac{1}{n}} - 1

Future Value
Year Deposit Amount Compounding Factor @12% Future Value
1 0 1.405 0.00
2 85000 1.254 106624.00
3 128600 1.120 144032.00
4 96780 1.000 96780.00
Future Value 347436.00

Present Value of Cash Outflow = \frac{260,000}{1+0.12}

Present Value of Cash Outflow = 232,142.86

MIRR = \frac{347,436}{232,142.86}^{\frac{1}{4}} - 1

MIRR = 10.61%

Therefore, Do Not Take Project because MIRR is less than Required rate of Return.

If you have any query related to question then feel free to ask me in a comment.Thanks.